if a & b are roots of a quadratic equation, please explain how
a^2+b^2 = a^2+2ab+b^2-2ab.
thank you.

Answer by Alan3354(69443)

a^2+b^2 = a^2+2ab+b^2-2ab.
------------
That's a^2 + b^2 = a^2 + b^2